Find the acceleration of a car that comes to a stop from a velocity of 60 m/s in a distance of 360 m.


1
Expert's answer
2022-03-02T14:38:58-0500

Using the kinematic formula (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas), obtain:


"v^2 = v_0^2-ad"

where "v=0,v_0=60m\/s" are the final and initial velocities respectively, "d = 360m" is the distance, and "a" is the acceleration (- sign since the car comes to stop). Thus, obtain:


"a = \dfrac{v_0^2}{d} = \dfrac{60^2}{360} = 10m/s^2"

Answer. 10m/s^2.